OpenStudy (anonymous):

if cos2X=1/3 and 0<=2X<=pi, find cosX

OpenStudy (anonymous):

2X = 70.53 degrees or 1.23 radians cos x = 0.8165

OpenStudy (anonymous):

i need it in terms of the unit circle?

OpenStudy (anonymous):

and an explanation would be great i havent done this in awile :/

OpenStudy (anonymous):

double angle results

OpenStudy (anonymous):

ok the angle 2x must be in the range 0 to pi or 0 to 180 as the cosine is positive there is only one result - that is in the first quadrant (0-90 degrees) so 2x = 70.53 and x = 70.53/ 2 then use calculator to find cos 70.53/2

OpenStudy (anonymous):

cos2x = cos^2-sin^2 = 2cos^2(x) -1

OpenStudy (anonymous):

there is the formula cos 2x = 2cos^2x - 1 which can be used as well

OpenStudy (anonymous):

therefore cos^2(x) = (1/2) ( 1+cos(2x) ) = (1/2) ( 1+ 1/3) = (1/2)(4/3) = (2/3) therefore cos(x) = +- sqrt(2/3)

OpenStudy (anonymous):

but 0<=2x<=pi so 0<=x<=pi/2 which means cos is in the first quadrant , so it is positive, so take the +case.

OpenStudy (anonymous):

\[\cos(x) = \sqrt{\frac{2}{3}}\]

OpenStudy (anonymous):

thank you so much